Driver of a car which is moving along a straight line road with a speed 90 km/h decides to decrease his speed to 36 km/h. If distance covered by the car during this period is 52.5 m, then acceleration


1
Expert's answer
2021-08-04T04:40:25-0400
"v^2=v_0^2+2as,""a=\dfrac{v^2-v_0^2}{2s},""a=\dfrac{(10\ \dfrac{m}{s})^2-(25\ \dfrac{m}{s})^2}{2\cdot52.5\ m}=-5\ \dfrac{m}{s^2}."